# Infinite families of irreducible polynomials over finite fields

### v1

@@ Abstract.
Let $p$ be a prime number and $q$ a power of $p$. Let ${\mathbb{F}}_{q}$ be the finite field with $q$ elements. For a positive integer $n$ and a polynomial $\varphi(X)\in{\mathbb{F}}_{q}[X]$, let $d_{n,\varphi}(X)$ denote the denominator of the $n$th iterate of $\frac{1}{\varphi(X)}$. The polynomial $\varphi(X)$ is said to be inversely stable over ${\mathbb{F}}_{q}$ if all polynomials $d_{n,\varphi}(X)$ are irreducible polynomial over ${\mathbb{F}}_{q}$ and distinct. In this paper, we characterize a class of inversely stable polynomials over ${\mathbb{F}}_{q}$. Actually, for $\varphi(X)=X^{p}-X+\xi\in{\mathbb{F}}_{q}[X]$, we provide a sufficient and necessary condition for $\varphi(X)$ to be inversely stable over ${\mathbb{F}}_{q}$.
@@
Finite fields, Irreducible polynomials, Iterations of rational functions.
@@ 1. Introduction
Constructing irreducible polynomials over finite fields is a central problem in finite field theory. Over the past few decades, numerous approaches have been developed to address this issue, particularly focusing on compositions and iterations of polynomials. One approach explores finding polynomials $f(X)$ and $g(X)$ such that $f(g(X))$ is an irreducible polynomial (see [2, 5, 6, 7, 16]). Another direction of research is identifying polynomials $h$ such that all of its iterates remain irreducible over a given finite field. Such polynomials are called stable over the finite field. Despite considerable attention to this topic (see [1, 8, 14, 15]), results remain sparse and the problem is far from being completely understood. For further discussions on different constructions of irreducible polynomials, one can refer to [3, 9, 10, 11, 12, 13] .
@@
Recently, the author [4] identified a class of inversely stable polynomials over the prime field ${\mathbb{F}}_{p}$, and proposed the following problem.
@@
In the case $e=t=1$, the author provided a complete solution to Problem 1.1 (see [4, Theorem 1.1]). However, for $e=t\geq 2$, it was shown that $f(X)=X^{p^{t}}+aX+b\in{\mathbb{F}}_{q}[X]$ is never inversely stable over ${\mathbb{F}}_{q}$. While a full solution to the general case remains elusive, this paper offers a partial solution for the case $e>t=1$.
@@
$\begin{cases}a_{3n}=2,c_{3n}=2w,d_{3n}=2w+2,a_{3n}c_{3n}^{-1}=w+2,{\rm Tr}_{\mathbb{F}_{9}}(a_{3n}c_{3n}^{-1})=2,\\ a_{3n+1}=2w+2,c_{3n+1}=2w+2,d_{3n+1}=2w+2,a_{3n}c_{3n}^{-1}=1,{\rm Tr}_{\mathbb{F}_{9}}(a_{3n+1}c_{3n+1}^{-1})=2,\\ a_{3n+2}=1,c_{3n+2}=2w,d_{3n+2}=1,a_{3n+2}c_{3n+2}^{-1}=2w+1,{\rm Tr}_{\mathbb{F}_{9}}(a_{3n+2}c_{3n+2}^{-1})=1.\end{cases}$
@@
The structure of the paper is as follows: In Section 2, we introduce a series of lemmas that lay the groundwork for the main results. Section 3 is devoted to proving Theorem 1.2.
@@ 2. Lemmas
[17, Equation (2.2)] Let $\alpha$ be an element of an extension field of $K$. Then the coefficient of the second highest term of the minimal polynomial of $\alpha$ over $K$ is equal to $-{\rm Tr}_{K(\alpha)/K}(\alpha)$.
@@
[17, Corollary 3.79] For $f(X)=X^{p}-X+\xi\in K[X]$, we have that $f(X)$ is irreducible over $K$ if and only if ${\rm Tr}_{K}(\xi)\neq 0$.
@@
Let $\alpha_{j}:=G^{(j)}(\infty)$ for each $j$ with $1\leq j\leq k$. Then $\alpha_{1},\ldots,\alpha_{k-1}$ are elements in $K$, and $g(\alpha_{k-1})=0$, implying that $g(X)$ has a root of $K$. However, by Lemma 2.4, $g(X)$ is irreducible over $K$ when ${\rm Tr}_{K}(\xi)\neq 0$, a contradiction. Therefore, $\infty\not\in(G^{(i)})^{-1}(\infty)$ for $i$. Now applying the result from Item (a), we see that
@@
Let $h_{1}(X)=X^{p}-X+\xi$, and for $2\leq j\leq n$, $h_{j}(X)=X^{p}-X+\frac{\xi\beta_{j-1}-1}{\beta_{j-1}}$. Clearly, $h_{i}(X)\in K_{i-1}[X]$ and $h_{i}(\beta_{i})=0$ for any $1\leq j\leq n$. It implies that $[K_{i}:K_{i-1}]\leq p$ for each $1\leq i\leq n$. So, we obtain that
@@ acknowledgment
This work was completed during the author’s academic visit to the Johann Radon Institute for Computational and Applied Mathematics (RICAM), Austrian Academy of Sciences. The author would like to express sincere gratitude to Professor Arne Winterhof at RICAM, for his careful reading of this paper and for providing valuable suggestions.

### v2

# The inverse stability of Artin-Schreier polynomials over finite fields

@@ Abstract.
Let $p$ be a prime number and $q$ a power of $p$. Let ${\mathbb{F}}_{q}$ be the finite field with $q$ elements. For a positive integer $n$ and a polynomial $\varphi(X)\in{\mathbb{F}}_{q}[X]$, let $d_{n,\varphi}(X)$ denote the denominator of the $n$th iterate of $\frac{1}{\varphi(X)}$. The polynomial $\varphi(X)$ is said to be inversely stable over ${\mathbb{F}}_{q}$ if all polynomials $d_{n,\varphi}(X)$ are irreducible polynomial over ${\mathbb{F}}_{q}$ and distinct. In this paper, we characterize a class of inversely stable polynomials over ${\mathbb{F}}_{q}$. More precisely, for $\varphi(X)=X^{p^{t}}+aX+b\in{\mathbb{F}}_{q}[X]$ with $t$ being a positive integer, we provide a sufficient and necessary condition for $\varphi(X)$ to be inversely stable over ${\mathbb{F}}_{q}$.
@@
Finite fields, inverse stability, Artin-Schreier polynomials.
@@ 1. Introduction
Constructing irreducible polynomials over finite fields is a central problem in finite field theory. Over the past few decades, numerous approaches have been developed to address this issue, particularly focusing on compositions and iterations of polynomials. One approach explores finding polynomials $f(X)$ and $g(X)$ such that $f(g(X))$ is an irreducible polynomial (see [4, 7, 8, 9, 19]). Another direction of research is identifying polynomials $h$ such that all of its iterates remain irreducible over a given finite field. Such polynomials are called stable over the finite field. Despite considerable attention to this topic (see [2, 3, 11, 17, 18]), results remain sparse and the problem is far from being completely understood. For further discussions on different constructions of irreducible polynomials, one can refer to [5, 10, 12, 13, 14, 15, 16] .
@@
The Artin-Schreier polynomials have the form $X^{p^{t}}+aX+b$, where $t$ is a positive integer and $a,b\in{\mathbb{F}}_{q}$. Recently, the author [6] identified a class of inversely stable Artin-Schreier polynomials over the prime field ${\mathbb{F}}_{p}$ and posed the following problem.
@@
In the case $e=t=1$, the author provided a complete solution to Problem 1.1 (see [6, Theorem 1.1]). However, for $e=t\geq 2$, it was shown that $f(X)=X^{p^{t}}+aX+b\in{\mathbb{F}}_{q}[X]$ is never inversely stable over ${\mathbb{F}}_{q}$. In this paper, we shall provide a solution to Problem 1.1 for the case $e>t=1$.
@@
$\begin{cases}a_{3n}=2,c_{3n}=2w,d_{3n}=2w+2,a_{3n}c_{3n}^{-1}=w+2,{\rm Tr}_{\mathbb{F}_{9}}(a_{3n}c_{3n}^{-1})=2,\\ a_{3n+1}=2w+2,c_{3n+1}=2w+2,d_{3n+1}=2w+2,a_{3n+1}c_{3n+1}^{-1}=1,{\rm Tr}_{\mathbb{F}_{9}}(a_{3n+1}c_{3n+1}^{-1})=2,\\ a_{3n+2}=1,c_{3n+2}=2w,d_{3n+2}=1,a_{3n+2}c_{3n+2}^{-1}=2w+1,{\rm Tr}_{\mathbb{F}_{9}}(a_{3n+2}c_{3n+2}^{-1})=1.\end{cases}$
@@
Theorem 1.2 offers a solution to Problem 1.1 for the case $e>t=1$. This case is central, as solutions to all other cases of Problem 1.1 can be obtained using the same method as that for the case $e>t=1$ (see Section 4). Consequently, Problem 1.1 is fully resolved in this sense.
The paper is organized as follows: Section 2 presents a series of lemmas that establish the foundation for the main results. Section 3 focuses on proving Theorem 1.2. Section 4 provides some remarks.
@@ 2. Lemmas
[20, Equation (2.2)] Let $\alpha$ be an element of an extension field of $K$. Then the coefficient of the second highest term of the minimal polynomial of $\alpha$ over $K$ is equal to $-{\rm Tr}_{K(\alpha)/K}(\alpha)$.
@@
[20, Corollary 3.79] For $f(X)=X^{p}-X+\xi\in K[X]$, we have that $f(X)$ is irreducible over $K$ if and only if ${\rm Tr}_{K}(\xi)\neq 0$.
@@
Let $\alpha_{j}:=G^{(j)}(\infty)$ for each $j$ with $1\leq j\leq k$. Then $\alpha_{1},\ldots,\alpha_{k-1}$ are elements in $K$, and $g(\alpha_{k-1})=0$, implying that $g(X)$ has a root of $K$. However, by Lemma 2.4, $g(X)$ is irreducible over $K$ when ${\rm Tr}_{K}(\xi)\neq 0$, a contradiction. Therefore, $\infty\not\in(G^{(i)})^{-1}(\infty)$ for every $i$ with $1\leq i\leq n$. Now applying the result from Item (a), we see that
@@
Let $h_{1}(X)=X^{p}-X+\xi$, and for $2\leq j\leq n$, $h_{j}(X)=X^{p}-X+\frac{\xi\beta_{j-1}-1}{\beta_{j-1}}$. Clearly, $h_{i}(X)\in K_{i-1}[X]$ and $h_{i}(\beta_{i})=0$ for any $1\leq i\leq n$. It implies that $[K_{i}:K_{i-1}]\leq p$ for each $1\leq i\leq n$. So, we obtain that
@@
4. Conclusion and Remarks
In this paper, we provided a complete characterization of the inverse stability of Artin–Schreier polynomials of the form $f(X)=X^{p^{t}}+aX+b\in\mathbb{F}_{q}[X]$ for the essential case $t=1$, $a=-1$, as posed in Problem 1.1. Our main result, Theorem 1.2, establishes that the inverse stability of $g(X)=X^{p}-X+b\in\mathbb{F}_{q}[X]$ is equivalent to the sequence-based condition $\operatorname{Tr}_{{\mathbb{F}}_{q}}(a_{n}c_{n}^{-1})\neq 0$ for all $n\geq 1$. This criterion enables us to directly verify the inverse stability of $g(X)$ by examining the absolute traces of specific sequences derived from the polynomial $g(X)$.
To characterize the inverse stability of $f(X)=X^{p^{t}}+aX+b$, we require $t\geq 1$ and $a\neq 0$. When $a=0$, $f(X)$ is reducible, as any $p$-polynomial over ${\mathbb{F}}_{q}$ with no nonzero roots in ${\mathbb{F}}_{q}$ permutes ${\mathbb{F}}_{q}$. Here, a $p$-polynomial has the form $\sum_{i}c_{i}X^{p^{i}}\in{\mathbb{F}}_{q}[X]$. In this paper, we only address Problem 1.1 in the case of $t=1$ and $a=-1$, but this case is fundamental. In fact, a complete solution to Problem 1.1 can be derived using the methods developed in this paper and the following observations, without repeating the proof process:
  1. (1)
    For $t=1$ and $a\neq 0$, a well-known result (see, for example, [21, Theorem 10.12]) states that $X^{p}+aX+b$ is irreducible over ${\mathbb{F}}_{q}$ if and only if there exists $a_{0}\in{\mathbb{F}}_{q}^{*}$ such that $a=-a_{0}^{p-1}$ and $\operatorname{Tr}(b/a_{0}^{p})\neq 0$. By applying the same methods as in this paper, a characterization similar to Theorem 1.2 can be obtained.
  2. (2)
    For $t\geq 2$ and $a\neq 0$, a result by Agou [1] indicates that a necessary condition for $f(X)$ to be irreducible over ${\mathbb{F}}_{q}$ is $t=p=2$. Thus, it suffices to characterize the inverse stability of $X^{4}+aX+b$ over $\mathbb{F}_{2^{e}}$. Moreover, according to Agou’s result, $X^{4}+aX+b$ is irreducible over $\mathbb{F}_{2^{e}}$ if and only if $e$ is odd, $a=a_{0}^{3}$, and $\operatorname{Tr}(b/a_{0}^{2})\neq 0$ for some $a_{0}\in\mathbb{F}_{2^{e}}^{*}$. Hence, the methods in this paper can be used to characterize the inverse stability of $X^{4}+aX+b$ over $\mathbb{F}_{2^{e}}$.
Finally, we note that when the sequences $\{a_{n}\}_{n\geq 1}$ and $\{c_{n}\}_{n\geq 1}$ are periodic, as shown in Example 1.4, the condition $\operatorname{Tr}_{{\mathbb{F}}_{q}}(a_{n}c_{n}^{-1})\neq 0$ can be checked by computing the trace for a finite number of terms within one period, rendering the criterion practical. In contrast, for non-periodic sequences, a key difficulty emerges: verifying $\operatorname{Tr}_{{\mathbb{F}}_{q}}(a_{n}c_{n}^{-1})\neq 0$ for increasingly large $n$ does not guarantee inverse stability, as the condition must hold for all $n$. This requires examining ever-larger $n$ without a finite stopping point, presenting a significant practical challenge.
@@ acknowledgment
The author expresses gratitude to the anonymous referees for their insightful and valuable comments, which significantly enhanced the paper’s presentation.
This work was conducted during the author’s academic visit to RICAM, Austrian Academy of Sciences. The author sincerely thanks Professor Arne Winterhof at RICAM for his thorough review of the manuscript and for offering invaluable suggestions.
This research was partially supported by the China Scholarship Council Fund (Grant No. 202301010002) and the Scientific Research Innovation Team Project of China West Normal University (Grant No. KCXTD2024-7).